# The Bathtub of European AI Governance: Identifying Technical Sandboxes as the Micro-Foundation of Regulatory Learning

### v1

@@ 6. Conclusion
The ambition of the EU AI Act to function as a future-proof, adaptive framework hinges on its ability to learn. However, the legal mechanisms for regulatory learning, such as reviewing cycles and standardisation requests, are insufficient without a robust technical foundation. By applying the bathtub model of social learning theory to the AI Act’s governance structure, we have identified a critical gap in the information flow, namely the translation of micro-level technical behaviour into macro-level regulatory signals. We proposed the AITS as the missing micro-foundation required to bridge this gap. Unlike the broader AIRS, the AITS provides the controlled experimentation environment necessary to reproducibly generate technical evidence, accounting for the deeply socio-technical nature of AI systems and their risks.
However, for this evidence to drive learning at scale, the AITS cannot function as a black box. We outlined requirements for machine-readable solutions, including a DSL, a standard vocabulary for metrics, and standardised assessment tool reporting to enable automated processing. Without semantic interoperability, the meso level, a complex network of actors, risks remaining the bottleneck in the regulatory learning space through inefficient data aggregation. Nevertheless, socio-political challenges, such as regulatory capture by powerful industry players and legislative inertia, remain and the proposed AITS, as a technical framework, cannot address them.
Ultimately, the success of the AI Act, in its ambition to balance governance with innovation, depends on operationalising this socio-technical infrastructure. If implemented correctly, the AITS can transform the compliance burden into valuable regulatory insight, both for companies and regulators. For future work, we intend to implement the components elaborated on in Section 4.

### v2

@@ 6. Conclusion
The ambition of the EU AI Act to function as a future-proof, adaptive framework hinges on its ability to learn. However, the legal mechanisms for regulatory learning, such as reviewing cycles and standardisation requests, are insufficient without a robust technical foundation. By applying the bathtub model of social learning theory to the AI Act’s governance structure, we have identified a critical gap in the information flow, namely the translation of micro-level technical behaviour into macro-level regulatory signals. We proposed AITSes as the missing micro-foundation required to bridge this gap. Unlike the broader AIRSes, AITSes provide the controlled experimentation environment necessary to reproducibly generate technical evidence, accounting for the deeply socio-technical nature of AI systems and their risks.
However, for this evidence to drive learning at scale, AITSes cannot function as black boxes. We outlined requirements for machine-readable solutions, including a DSL, a reference ontology for metrics, and standardised assessment tool reporting to enable automated processing. Without semantic interoperability, the meso level, characterised by a complex network of actors, risks remaining the bottleneck in the regulatory learning space due to inefficient data aggregation. Nevertheless, broader socio-political challenges, including risks of regulatory capture by well-resourced industry actors and forms of legislative inertia, remain outside the scope of technical solutions. As a result, AITSes, while enabling structured evidence generation and regulatory learning, cannot by themselves address these challenges.
While this work has focused exclusively on the AI Act, including the mapping of its stakeholders and its pathways for regulatory learning, the proposed AITS characteristics are not intrinsically tied to the EU context. Similar analyses could be conducted for other geographic and regulatory settings. The technical properties required to enable structured evidence generation and regulatory learning are sufficiently generic to support such extensions. However, it is within the EU that these mechanisms are expected to have the greatest impact, given the diversity of implementation practices across Member States and the resulting fragmentation in how technical evidence is generated, shared, and aggregated. Ultimately, the success of the AI Act, in its ambition to balance governance with innovation, depends on operationalising this socio-technical infrastructure. If implemented correctly, AITSes can transform compliance efforts into valuable regulatory insight for both providers and regulators.
For future work, we intend to implement the components of an AITS, as outlined in Section 4, to enable structured evidence generation, assess their regulatory impact in practice, and compare their application with emerging approaches in other geographic contexts.
Acknowledgement
This work is supported by the Luxembourg AI Factory.